OpenStudy (anonymous):

k = am+3mx what is the value of m in terms of a,k and x

OpenStudy (luigi0210):

If you mean solve for m.. then: \[\LARGE k=am+3mx\] Factor out m: \[\LARGE k=m(a+3x)\] Now divide: \[\LARGE \frac{k}{a+3x}=m\]

OpenStudy (anonymous):

you can start on the right side of the equation by noticing that both terms have "m" in them... "am" and "3mx" So you can factor the right side like: m(a + 3x)

OpenStudy (anonymous):

and welcome 2 os ;)